# TOB-SVD: Total-Order Broadcast with Single-Vote Decisions in the Sleepy Model

### v2

@@ Abstract
Over the past years, distributed consensus research has extended its focus towards addressing challenges in large-scale, permissionless systems, such as blockchains. This shift is characterized by the need to accommodate dynamic participation, contrasting the traditional approach of a static set of continuously online participants. Works like Bitcoin and the sleepy model have set the stage for this evolving framework.
Notable contributions from Momose and Ren (CCS 2022) and subsequent works have introduced Total-Order Broadcast protocols leveraging Graded Agreement primitives and supporting dynamic participation. However, these approaches often require multiple phases of voting per decision, creating a potential bottleneck for real-world large-scale systems.
Addressing this, our paper introduces TOB-SVD, a novel Total-Order Broadcast protocol in the sleepy model, which is resilient to up to 1/2 of adversarial participants. TOB-SVD requires only a single phase of voting per decision in the best case and achieves lower expected latency compared to existing approaches offering the same optimal adversarial resilience. This work paves the way to more practical Total-Order Broadcast protocols to be implemented in real-world systems where a large number of participants are involved simultaneously and their participation level might fluctuate over time.
@@ 1 Introduction
Distributed consensus research has expanded in recent years to address the challenges posed by large-scale, permissionless systems, such as blockchains. In contrast to traditional methods for solving consensus, which assume participants to be continuously online and actively contributing to the protocol, these new consensus protocols aim at accommodating dynamicity in the participation level among network participants [11]. We refer to consensus (or, in our context, Total-Order Broadcast) protocols supporting dynamic participation as dynamically available.
@@
Momose and Ren’s research [10] laid the foundation for dynamically available Total-Order Broadcast (TOB) protocols with deterministic Safety, sparking a series of subsequent research [8, 9, 7, 3]. Notably, the protocols stemming from these works share a common structural theme: they all leverage a Graded Agreement (GA) primitive [^1] In Graded Agreement, a grade is associated with any decision where the grade intuitively indicates how strong the agreement is., albeit with different implementations and properties. For example, the Total-Order Broadcast protocol proposed by Momose and Ren [10] (MR) utilizes a Graded Agreement protocol resilient to 1/2 adversarial participants. However, their TOB protocol is rather impractical due to high latency. In contrast, the subsequent work by Malkhi, Momose, and Ren [8] (MMR) proposes two protocols (1/3MMR and 1/4MMR) which improve latency but lower the adversarial tolerance. A later refinement by the same authors [9] (MMR2) reverts to tolerating minority corruption, while maintaining comparable latency. A concurrent and independent work [7] (GL) also uses a Graded Agreement primitive (called Commit-Adopt in their work) [^2] To be precise, GL employs a primitive called Adopt-Commit, which slightly differs from the Graded Agreement (GA) primitive [2]. For simplicity, we will refer to it as GA in this work, although it is important to note that there are differences between the two primitives..
@@
The three currently existing dynamically available protocols with optimal adversarial resilience and latency constant in the security parameter, i.e., [10, 9, 7], all share the drawback of requiring multiple phases of voting [^3] A voting phase is a point in time when participants perform computations based on received messages and subsequently send messages whose content depends on these computations. for each new block (i.e., decision). In practice, these protocols all operate in views, and within each view a block is proposed and a decision is made. To do so, two [7], three [9], and five [10] instances of Graded Agreement are invoked within a view, each including four, three, and two phases of voting, respectively. This poses a challenge to their practical application in real-world systems, particularly in scenarios involving a multitude of simultaneous participants, such as those found in most of existing blockchains. In fact, a key factor for the practical application in such systems is reducing the number of voting phases per new block as much as possible. This is because of various reasons. First, due to the large number of participants involved, voting phases are computationally intensive which results in an extra computation delay added on top of the network latency. Additionally, these protocols often require a signature aggregation process where messages are first sent to aggregators who then distribute the aggregated signatures, causing voting phases to require double the normal network latency. For instance, the Ethereum’s consensus protocol divides its slot – the unit of time in which a new block is proposed and voted – in three phases taking up $\Delta=4s$ each. One $\Delta$ is for block propagation and two $\Delta$ are for vote propagation, one of which is reserved to propagation and aggregation of votes, and one to propagation of aggregated votes. Effectively, a voting phase in Ethereum takes $2\Delta$[1].
@@
    Introduction of a dynamically available Total-Order Broadcast protocol that tolerates up to 1/2 adversarial participants.
@@
The remainder of this work is structured as follows. We present and compare related work in Section 2. Section 3 presents the system model and essential definitions. Importantly, we revisit the sleepy model [11], expanding upon the notation first introduced by Malkhi, Momose, and Ren [9]. Section 4 gives a technical outline of our results presenting two GA protocols, one with two grades which serves as a foundation for the subsequent GA protocol with three-grades. In Section 5, we analyze both Graded Agreement protocols. The Graded Agreement protocol with three-grades is then used in our Total-Order Broadcast protocol presented in Section 6. Finally, conclusions are drawn in Section 7.
@@ 2 Related Work
Voting phases per new block in the best case $1$ $10$ $3$ $5$ $2$ $1$
Voting phases per new block in the expected case $2$ $20$ $12$ $10$ $4$ $2$
Table 1: Comparison of dynamically available TOB protocols with deterministic Safety.
To provide a clearer motivation for our results, we first compare the outcomes of our work with those of other protocols. To facilitate this comparison, we introduce key metrics that will aid in evaluating the performance and characteristics of different protocols.
We assume that upon submission, transactions are immediately added to a transaction pool from which participants can retrieve and validate them using a specified validity predicate [^4] For simplicity, we assume that the validity of a transaction is evaluated independently of any other transaction already included in the log. A transaction is valid according to a global, efficiently computable predicate $P$, known to all validators. The specific details of this predicate are omitted. before batching them into blocks. These blocks are then appended to an existing sequence of other blocks, or a log (Section 3.2), and such a resulting log is proposed for decision by the consensus protocol. We assume that honest validators batch into any proposed block any valid transaction included in the transaction pool that is not already included in the log that the proposed block is appended to.
The confirmation time of a transaction is defined as the time between its submission (from users) and the decision regarding the log containing that transaction.
We refer to best-case latency as the minimum possible confirmation time of a transaction, where the minimum is taken over all possible submission times. Simply put, it represents the shortest time between a proposal and its decision.
@@
We compare all existing dynamically available Total-Order Broadcast protocols that ensure deterministic Safety. Note that Fig. 4 in Table 1 highlight our contributions, which are presented in detail in Section 6.
@@
The expected latency of our protocol is $10\Delta$, which improves on prior works: $14\Delta$, $20\Delta$, and $32\Delta$ for the MMR2, GL, and MR protocols, respectively. Specifically, under our assumption regarding leader election (Section 3.3), our protocol utilizes a VRF-based leader election mechanism. Each transaction is expected to be batched by an honest proposer within two views, resulting in an expected confirmation time of $10\Delta$. In contrast, the 1/3MMR and 1/4MMR protocols must reduce their adversary tolerance to achieve comparable latency improvements.
@@
In the best case, our protocol requires only one voting phase per new block, which improves on prior works: 10, 5, 3, and 2 voting phases per new block for the MR, GL, MMR2, GL, and 1/3MMR protocols, respectively. Like our protocol, the 1/4MMR protocols only needs a single voting phase.
Analyzing the expected case, the GL protocol uses 10 voting phases. The MMR2 protocol requires 4 GA instances, each involving 3 voting phases, resulting in a total of 12 voting phases per new block, while the MR protocol requires 10 GA instances with 2 voting phases each, totaling 20 voting phases per new block. Similarly, the 1/3MMR and 1/4MMR protocols needs 4 and 2 voting phases, respectively. Finally, our protocol needs only 2 voting phases per new block.
Our protocol achieves the same expected communication complexity as MMR2, GL, and MR which is $\mathit{O}(Ln^{3})$, where $L$ represents the block size. Both 1/3MMR and 1/4MMR have a communication complexity of only $O(Ln^{2})$ as they do not forward received messages, but, as highlighted above, this goes to the detriment of adversarial resiliency.
Finally, our protocol requires a stabilization period as a fundamental assumption for its security. Other protocols, such as 1/3MMR, 1/4MMR, and MMR2, do not require such a period. However, in practice, both these protocols and ours necessitate of a recovery mechanism to allow nodes that wake up to recover messages that were sent to them while asleep, since assuming that messages are buffered and delivered immediately is not very practical.
Such recovery protocol typically requires that, upon waking up, a node sends a $\mathsf{RECOVERY}$ message to other nodes. These nodes then send back any messages that the newly awakened node may have missed while asleep and that could impact future decisions. The node that wakes up is required to remain awake until it receives responses to the $\mathsf{RECOVERY}$ messages it has sent out. Only then is it considered awake in the context of the adversary model constraints. Such a period is, in practice, at least $2\Delta$.
Therefore, in practice, we share with other related works the requirement for a period of stable honest node participation. For our protocol, this requirement is stronger than that of 1/3MMR, 1/4MMR, and MMR2, as we need stable participation during both the recovery period and the stabilization period $T_{s}$. This means that, overall, our protocol’s stabilization assumption is $T_{s}$ longer than that of other protocols. However, note that, in practice, when a node wakes up, it may need to download a significant amount of data to catch up, potentially requiring more than just $2\Delta$. For example, if the recovery procedure takes several hours, assuming an additional stabilization period of $T_{s}=2\Delta$ should not significantly impact the overall protocol.
@@ 3.1 System model
Validators. We consider a system of $n$ validators $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ in a message-passing system with an underlying peer-to-peer dissemination protocol (e.g., a gossip protocol) over a synchronous network with delay bound $\Delta>0$. We assume that validators have synchronized clocks. Each validator is uniquely identified by a cryptographic identity, and their public keys are common knowledge. We use the notation ${\langle m\rangle}_{i}$ to indicate that a message $\langle m\rangle$ is signed by validator $v_{i}$. A protocol for $\mathcal{V}$ consists of a collection of programs with instructions for all validators.
Failures. A validator that follows its protocol during an execution is called honest. At any time, honest validators forward any message received. On the other hand, a faulty validator may deviate arbitrarily from its specification, e.g., when corrupted by an adversary; we refer to such validators as Byzantine.
Adversary and sleepy model. The adversary can permanently corrupt honest validators throughout the execution in a mildly adaptive way. To be more specific, if the adversary corrupts an honest validator $v_{i}$ at time $t$, then $v_{i}$ becomes Byzantine only at time $t+\Delta$. In other words, we assume that the adversary does not have the capability to corrupt honest validators immediately. Instead, there is a delay, represented by $\Delta$, before it can do so. This delay is essential for the effective operation of a VRF-based leader election, as detailed in Section 3.2.
The adversary can also fully adaptively either put validators to sleep, preventing them from participating in the protocol, or wake them up, enabling them to participate in the protocol. Validators subjected to the sleeping state are referred to as asleep, while those actively participating in the protocol are designated awake. We assume that, upon waking up, validators immediately receive all messages they should have received while asleep. This is in line with other works [11, 10, 8, 9]. In other words, we are considering (a variant of) the sleepy model [11], which we now fully specify.
@@
To be precise, for any time $t\geq 0$, let $H_{t}$ be the set of awake honest validators at time $t$, and let $B_{t}$ be the set of Byzantine validators at time $t$ [^6] If $t<0$ we define $H_{t}:=\mathcal{V}$ and $B_{t}:=\emptyset$.. We assume that $t_{1}\leq t_{2}\implies B_{t_{1}}\subseteq B_{t_{2}}$, i.e., we assume $B_{t}$ to be monotonically non-decreasing [^7] The reason for this assumption is that Byzantine validators can always transfer their private keys to other Byzantine validators prior to entering a sleep state, enabling these keys to be used indefinitely.. For the interval $[t_{1},t_{2}]$ and a time $t_{3}$, we define $H_{t_{1},t_{2},t_{3}}=(\bigcap_{t\in[t_{1},t_{2}]}H_{t})\setminus{B_{t_{3}}}$ to be the set containing the honest validators that were awake between $t_{1}$ and $t_{2}$ and remain honest up to time $t_{3}$.
Let $T_{b}$ and $T_{s}$ be non-negative integers representing two specific time constants. Furthermore, let $\rho\geq 1$ denote a predetermined failure ratio of honest to Byzantine validators. A system is compliant with the $(T_{b},T_{s},\rho)$-sleepy model if and only if, for every time instance $t\geq 0$, the following condition is satisfied:
$|H_{t-T_{s},t,t+T_{b}}|>\rho\cdot|B_{t+T_{b}}|$ (1)
In other words, the validators that are awake and honest throughout the period $[t-T_{s},t]$ and still not corrupted by time $t+T_{b}$ outnumber the Byzantine validators at $t+T_{b}$ by a factor of $\rho$.
The parameter $T_{b}$ represents the duration for which backward simulation is considered. In other words, Byzantine validators are counted for an extra time $T_{b}$ backward with respect to $t$. Additionally, $T_{b}$ captures the requirement that not too many honest validators become corrupted within $T_{b}$ after performing a specific action, such as voting, in order to prevent them from sending contradictory messages shortly afterward. Our model is additionally augmented with $T_{s}$ in order to encapsulate a stable participation requirement. This is emphasized by focusing on honest validators who remain active throughout the time span $T_{s}$.

### v3

@@ Abstract
Over the past years, distributed consensus research has expanded its focus to address challenges in large-scale, permissionless systems, such as blockchains. This shift reflects the need to accommodate dynamic participation, in contrast to the traditional model of a static set of continuously online validators. Works like Bitcoin and the sleepy model have laid the groundwork for this evolving framework.
Notable contributions by Momose and Ren (CCS 2022), along with subsequent research, have introduced Total-Order Broadcast protocols that leverage Graded Agreement primitives and support dynamic participation. However, these approaches often require multiple phases of voting per decision, which can create a bottleneck for real-world, large-scale systems.
To address this, our paper introduces TOB-SVD, a novel Total-Order Broadcast protocol in the sleepy model that is resilient to up to 1/2 Byzantine validators. TOB-SVD requires only a single phase of voting per decision in the best case and achieves lower expected latency compared to existing approaches offering the same optimal adversarial resilience. This work paves the way for more practical Total-Order Broadcast protocols that can be implemented in real-world systems involving a large number of validators with fluctuating participation over time.
@@ 1 Introduction
Distributed consensus research has expanded in recent years to address the challenges posed by large-scale, permissionless systems, such as blockchains. In contrast to traditional methods for solving consensus, which assume validators to be continuously online and actively contributing to the protocol, these new consensus protocols aim at accommodating dynamicity in the participation level among network validators [14]. We refer to consensus (or, in our context, Total-Order Broadcast) protocols supporting dynamic participation as dynamically available.
@@
Momose and Ren’s research [12] laid the foundation for dynamically available Total-Order Broadcast (TOB) protocols with deterministic Safety, sparking a series of subsequent research [10, 11, 9, 4]. Notably, the protocols stemming from these works share a common structural theme: they all leverage a Graded Agreement (GA) primitive [^1] In Graded Agreement, a grade is associated with any decision where the grade intuitively indicates how strong the agreement is. [2], albeit with different implementations and properties. For example, the Total-Order Broadcast protocol proposed by Momose and Ren [12] (MR) utilizes a Graded Agreement protocol resilient to 1/2 Byzantine validators. However, their TOB protocol is rather impractical due to high latency. In contrast, the subsequent work by Malkhi, Momose, and Ren [10] (MMR) proposes two protocols (1/3MMR and 1/4MMR) which improve latency but lower the adversarial tolerance. A later refinement by the same authors [11] (MMR2) reverts to tolerating minority corruption, while maintaining comparable latency. A concurrent and independent work [9] (GL) also uses a Graded Agreement primitive (called Commit-Adopt in their work). [^2] To be precise, GL employs a primitive called Commit-Adopt, which slightly differs from the Graded Agreement (GA) primitive [2]. For simplicity, we will refer to it as GA in this work, although it is important to note that there are differences between the two primitives.
@@
The three currently existing dynamically available protocols with optimal adversarial resilience and latency constant in the security parameter, i.e., [12, 11, 9], all share the drawback of requiring multiple phases of voting [^3] A voting phase is a point in time when every honest validator performs computations based on the messages they have received and subsequently sends a new message. Here, “new” refers to a message that has never been sent before, with its content determined by these computations. for each new block (i.e., decision). In practice, these protocols all operate in views, and within each view a block is proposed and a decision is made. To do so, two [9], three [11], and five [12] instances of Graded Agreement are invoked within a view, each including four, three, and two phases of voting, respectively. This poses a challenge to their practical application in real-world systems, particularly in scenarios involving a multitude of simultaneous validators, such as those found in most of existing blockchains. In fact, a key factor for the practical application in such systems is reducing the number of voting phases per new block as much as possible. This is because of various reasons. First, due to the large number of validators involved, voting phases are computationally intensive which results in an extra computation delay added on top of the network latency. Additionally, these protocols often require a signature aggregation process where messages are first sent to aggregators who then distribute the aggregated signatures, causing voting phases to require double the normal network latency. For instance, the Ethereum’s consensus protocol divides its slot – the unit of time in which a new block is proposed and voted – in three phases taking up $\Delta=4s$ each. One $\Delta$ is for block propagation and two $\Delta$ are for vote propagation, one of which is reserved to propagation and aggregation of votes, and one to propagation of aggregated votes. Effectively, a voting phase in Ethereum takes $2\Delta$[1].
@@
    Introduction of a dynamically available Total-Order Broadcast protocol that tolerates up to 1/2 Byzantine validators.
@@
Expanding on point 3, a recent research direction demonstrates how combining a dynamically available total-order broadcast protocol with a finality gadget – a partially synchronous total-order broadcast protocol – can produce a secure ebb-and-flow protocol [13]. Specifically, an ebb-and-flow protocol consists of two components: (1) a dynamically available total-order broadcast protocol that guarantees safety and liveness under synchrony ($\mathsf{GST}=0$) at all times, and (2) a finalizing protocol that ensures safety at all times and liveness after $\max(\mathsf{GST},\mathsf{GAT})$, where $\mathsf{GAT}$ represents a global awake time after which all honest validators come online.
Practical research has further explored [8], within the context of the Ethereum consensus protocol, how to combine a probabilistically safe dynamically available total-order broadcast protocol [6] and a finality gadget in order to achieve single-slot finality [3]. Building on this, we strongly believe that similar results can be achieved by replacing their dynamically available protocol with the protocol presented in this work, potentially reducing message sizes.
The remainder of this work is structured as follows. We present and compare related work in Section 2. Section 3 presents the system model and essential definitions. Importantly, we revisit the sleepy model [14], expanding upon the notation first introduced by Malkhi, Momose, and Ren [11]. Section 4 recalls a foundational result that serves as the starting point for our protocol. Section 5 details our contributions. Specifically, Section 5.1 and Section 5.2 introduce two Graded Agreement protocols: the first, a two-grade GA protocol, lays the groundwork for the more advanced three-grade GA protocol. Section 5.3 then presents TOB-SVD, our dynamically available Total-Order Broadcast protocol. We provide a comprehensive analysis of both Graded Agreement protocols and TOB-SVD in Section 6. Finally, Section 7 concludes the paper.
@@ 2 Related Work
Voting phases per new block in the best case 11 1010 33 55 22 11
Voting phases per new block in the expected case 22 2020 1212 1010 44 22
Communication complexity $O(Ln^{3})$ $O(Ln^{3})$ $O(Ln^{3})$ $O(Ln^{3})$ $O(Ln^{2})$ $O(Ln^{2})$
Table 1: Comparison of dynamically available TOB protocols with deterministic Safety, where $L$ represents the block size and $n$ refers to the number of validators.
To provide a clearer motivation for our results, we first compare the outcomes of our work with those of other protocols. To facilitate this comparison, we introduce key metrics to help evaluate the performance and characteristics of protocols.
In the following, we assume that upon submission, transactions are immediately added to a transaction pool from which validators can retrieve and validate them using a specified validity predicate [^4] For simplicity, we assume that the validity of a transaction is evaluated independently of any other transaction already included in the log. A transaction is valid according to a global, efficiently computable predicate $P$, known to all validators. The specific details of this predicate are omitted. before batching them into blocks. These blocks are then appended to an existing sequence of other blocks, or a log (Section 3.2), and such a resulting log is proposed for decision by the consensus protocol. We assume that honest validators batch into any proposed block any valid transaction included in the transaction pool that is not already included in the log that the proposed block is appended to.
The confirmation time of a transaction is defined as the time between its submission (from a user) and the decision regarding the log containing that transaction.
We refer to best-case latency as the minimum possible confirmation time of a transaction, where the minimum is taken over all possible submission times. In practice, this corresponds to the shortest time between a proposal and its decision.
@@
In Table 1 we compare all existing dynamically available Total-Order Broadcast protocols that ensure deterministic Safety, where $L$ represents the block size and $n$ refers to the number of validators.
@@
Our protocol’s expected latency is $10\Delta$, which improves on prior works: $14\Delta$, $20\Delta$, and $32\Delta$ for the MMR2, GL, and MR protocols, respectively. Specifically, under our assumption regarding leader election (Section 3.3), our protocol utilizes a Verifiable Random Function (VRF)-based leader election mechanism. Each transaction is expected to be batched by an honest proposer within two views, resulting in an expected confirmation time of $10\Delta$. In contrast, the 1/3MMR and 1/4MMR protocols reduce their adversary tolerance to achieve comparable latency improvements.
@@
In the best case, our protocol requires only one voting phase per new block, which improves on prior works: 10, 5, 3, and 2 voting phases per new block for the MR, GL, MMR2, and 1/3MMR protocols, respectively. Like our protocol, the 1/4MMR protocols only needs a single voting phase.
Analyzing the expected case, the GL protocol uses 10 voting phases. The MMR2 protocol requires 4 GA instances, each involving 3 voting phases, resulting in a total of 12 voting phases per new block, while the MR protocol requires 10 GA instances with 2 voting phases each, totaling 20 voting phases per new block. Similarly, the 1/3MMR and 1/4MMR protocols need 4 and 2 voting phases, respectively. Finally, our protocol needs only 2 voting phases per new block.
Our protocol achieves the same expected communication complexity as MMR2, GL, and MR which is $\mathit{O}(Ln^{3})$, where $L$ represents the block size. Both 1/3MMR and 1/4MMR have a communication complexity of only $O(Ln^{2})$ as they do not forward received messages, while all others do. However, as highlighted above, this goes to the detriment of adversarial resiliency.
Finally, our protocol requires a stabilization period of $2\Delta$ as a fundamental assumption for its security. Other protocols, such as 1/3MMR, 1/4MMR, and MMR2, do not require such a period. However, in practice, both these protocols and ours necessitate of a recovery mechanism to allow validators that wake up to recover messages that were sent to them while asleep, since assuming that messages are buffered and delivered immediately is not very practical.
Such recovery protocol typically requires that, upon waking up, a validator sends a $\mathsf{RECOVERY}$ message to other validators. These validators then send back any messages that the newly awakened validator may have missed while asleep and that could impact future decisions. The validator that wakes up is required to remain awake until it receives responses to the $\mathsf{RECOVERY}$ messages it has sent out. Only then it is considered awake in the context of the adversary model constraints. Such a period is, in practice, at least $2\Delta$.
Therefore, in practice, we share with other related works the requirement for a period of stable honest validator participation. For our protocol, this requirement is stronger than that of 1/3MMR, 1/4MMR, and MMR2, as we need stable participation during both the recovery period and the stabilization period. This means that, overall, our protocol’s stabilization assumption is $2\Delta$ longer than that of other protocols. However, note that, in practice, when a validator wakes up, it may need to download a significant amount of data to catch up, potentially requiring more than just $2\Delta$. For example, if the recovery procedure takes several hours, assuming an additional stabilization period of $2\Delta$ should not significantly impact the overall protocol.
@@ 3.1 System model
Validators. We consider a system of $n$ validators $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ in a message-passing system with an underlying synchronous network with delay bound $\Delta>0$. We assume that validators have synchronized clocks. Each validator is uniquely identified by a cryptographic identity, and their public keys are common knowledge. We use the notation ${\langle m\rangle}_{i}$ to indicate that a message $\langle m\rangle$ is signed by validator $v_{i}$. A protocol for $\mathcal{V}$ consists of a collection of programs with instructions for all validators.
Honest and Byzantine validators. A validator that follows its protocol during an execution is called honest. At any time, honest validators forward any message received. On the other hand, a faulty validator may deviate arbitrarily from its specification, e.g., when corrupted by an adversary; we refer to such validators as Byzantine.
Adversary and sleepy model. We assume an adversary capable of permanently corrupting honest validators throughout the execution in a mildly adaptive manner. To be more specific, if the adversary corrupts an honest validator $v_{i}$ at time $t$, then $v_{i}$ becomes Byzantine only at time $t+\Delta$. In other words, we assume that the adversary does not have the capability to corrupt honest validators immediately. Instead, there is a delay, represented by $\Delta$, before it can do so. This delay is essential for the effective operation of a VRF-based leader election, as detailed in Section 3.3.
The adversary can also fully adaptively either put validators to sleep, preventing them from participating in the protocol, or wake them up, enabling them to participate in the protocol. Validators subjected to the sleeping state are referred to as asleep, while those actively participating in the protocol are designated awake. We assume that, upon waking up, validators immediately receive all messages they should have received while asleep. [^6] This assumption, while not practical for real-world systems, serves as a theoretical framework for analyzing the protocol, consistent with prior works [14, 12, 10, 11]. As briefly discussed in Section 2, incorporating a recovery procedure into the protocol could eliminate the need for this assumption, aligning it with practical considerations. However, the design and integration of such a recovery procedure in our protocol fall outside the scope of this work. In other words, we are considering (a variant of) the sleepy model [14], which we now fully specify.
@@
To be precise, for any time $t\geq 0$, let $H_{t}$ be the set of awake honest validators at time $t$, and let $B_{t}$ be the set of Byzantine validators at time tt [^7] If $t<0$ we define $H_{t}:=\mathcal{V}$ and $B_{t}:=\emptyset$.. We assume that $t_{1}\leq t_{2}\implies B_{t_{1}}\subseteq B_{t_{2}}$, i.e., we assume $B_{t}$ to be monotonically non-decreasing [^8] The reason for this assumption is due to forward simulation.. For the interval $t_{1},t_{2}]$, we define $H_{t_{1},t_{2}}=(\bigcap_{t\in[t_{1},t_{2}]}H_{t})$ to be the set containing the honest validators that were awake between $t_{1}$ and $t_{2}$.
Let $T_{b}$ and $T_{s}$ be non-negative integers representing two specific time constants. The parameter $T_{b}$ represents the duration for which backward simulation [11] is considered. In other words, Byzantine validators are counted for an extra time $T_{b}$ backward. In other words, $T_{b}$ captures the requirement that not too many honest validators become corrupted within $T_{b}$ after performing a specific action, such as voting, in order to prevent them from sending contradictory messages shortly afterward. Our model is additionally augmented with the stabilization period $T_{s}$ in order to encapsulate a stable participation requirement. This is emphasized by focusing on honest validators who remain awake throughout the time span $T_{s}$.
Then, we define the active validators at time $t$ to be the set $H_{t-T_{s},t}\cup B_{t+T_{b}}$. Intuitively, this corresponds to the smallest set of validators that might [^9] By “might” we mean that there exists an execution where this happens. send a message during a GA instance [^10] This concept will be formally introduced later. starting at time $t$ and lasting $T_{b}$. More generally, this set corresponds to the smallest set of validators that might influence outputs/decisions dependent on messages sent by honest validators at time $t$.
Furthermore, let $\rho\leq\frac{1}{2}$ denote a predetermined failure ratio of Byzantine to active validators.
A system is compliant with the $(T_{b},T_{s},\rho)$-sleepy model if and only if, for every time $t\geq 0$, the following condition
is satisfied:
$|B_{t+T_{b}}|<\rho\cdot|H_{t-T_{s},t}\cup B_{t+T_{b}}|$ (1)